On January 1, 2015, Swift Corporation, a small manufacturer of machine tools, acquired new industrial equipment for $1,320,000. The new equipment had a useful life of five years and the residual value was estimated to be $60,000. Swift estimates that the new equipment can produce 14,500 machine tools in its first year. It estimates that production will decline by 1,000 units per year over the equipment’s remaining useful life.

The following depreciation methods may be used: (1) straight-line, (2) double-declining-balance, and (3) unit of production. For tax purposes, the CCA class is Class 10—30%.

Prepare a schedule showing the amount of accumulated depreciation at December 31, 2017, under straight-line method, double-declining balance method and units-of-output method.

Straight-line Double-declining balance Units-of-output
2015 $

$

$

2016 $

$

$

2017 $

$

$

Which of the three depreciation methods would maximize net income for financial statement reporting purposes for the three-year period ending December 31, 2017?

_________________________

would maximize net income for financial statement reporting purposes for the three year period ending December 31, 2017.
Answer #1

Solution:

Straight line method
Cost 1320000
Salvage value 60000
Depreciable value 1260000
Life 5
Annual depreciation 252000
Double -declining method
Year Depreciation expense
2015 (1320000/5)*2 = 528000
2016 (1320000-528000)*40% = 316800
2017 (1320000-844800)*40% = 190080
units of output Year Expected output Depreciation expense
Cost 1320000 2015 14500 292320
Salvage value 60000 2016 13500 272160
Depreciable value 1260000 2017 12500 252000
Expected Life output 62500 2018 11500
Unit depreciation expense 20.16 2019 10500
SLM Double declining balance Units-of-output
2015 $252000 $528000 $292320
2016 $252000 $316800 $272160
2017 $252000 $190080 $252000
Total $756000 $1034880 $816480

Answer : Straight line method would maximize net income for three year period ending December 2017.